# Ionic association and Wien effect in 2D confined electrolytes

### v1

@@ Conclusion
We expect that the effective potential approach used in this paper could be extended to explore other materialsKavokine, Robin, and Bocquet (2022), but also other geometries, for example the case of multiple ionic layersCoquinot et al. (2024), by changing the effective interaction potential accordingly.
Overall, we obtain excellent agreement between our analytical models and numerical results, for both the pairing transition and the 2nd Wien effect. In particular, we find that in the pair-dominated regime, the ionic current behaves like a power law of the applied field, with a non-universal exponent that can be predicted from analytical field theories. This regime, where the conductivity strongly vanishes at low electric field, can be considered as ionic coulomb blockade situation – although no gating dependence is considered here –.
@@
Finally, the emergence of strongly non-linear conduction effects in 2D is a richness which can be exploited to develop nanofluidic systems with advanced properties, such as memristors Robin et al. (2022); Robin and Bocquet (2023). This is an opportunity which will certainly result in further developments in this active domain Kamsma et al. (2024); Emmerich et al. (2024); Paulo et al. (2023)
@@ Acknowledgements
The authors thank B. Coquinot and G. Monet for fruitful discussions. LB acknowledges support from ERC-Synergy grant agreement No.101071937, n-AQUA. PR acknowledges support from the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No.101034413).

### v2

@@ Conclusion
We expect that the effective potential approach used in this paper could be extended to explore other materialsKavokine, Robin, and Bocquet (2022), but also other geometries, for example the case of multiple ionic layersCoquinot et al. (2024), by changing the effective interaction potential accordingly. Models of ionic clusters dynamics are also useful to study ionic liquids, where many parallels can be drawn for the relations between ion correlations and transport phenomena Feng et al. (2019).
Overall, we obtain excellent agreement between our analytical models and numerical results, for both the pairing transition and the 2nd Wien effect. In particular, we find that in the pair-dominated regime, the ionic current behaves like a power law of the applied field, with a non-universal exponent that can be predicted from analytical field theories. This regime, where the conductivity strongly vanishes at low electric field, can be considered as ionic coulomb blockade situation – although no gating dependence is considered here.
@@
Finally, the emergence of strongly non-linear conduction effects in 2D is a richness which can be exploited to develop nanofluidic systems with advanced properties, such as memristors Robin et al. (2023a); Robin and Bocquet (2023). This is an opportunity which will certainly result in further developments in this active domain Kamsma et al. (2024); Emmerich et al. (2024); Paulo et al. (2023)
IV Supplementary Material
Supplementary Material includes details of numerical simulations, including the values of all numerical parameters. We also detail there the procedures to accurately extract the average number of pairs from numerical correlation functions, as well as the IV curve exponent $\alpha$.
@@ Acknowledgements
The authors thank B. Coquinot and G. Monet for fruitful discussions. LB acknowledges support from ERC-Synergy grant agreement No.101071937, n-AQUA. PR acknowledges support from the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No.101034413.